\begin{thebibliography}{33}

%\bibitem{Adams}
%G. T. Adams, N. S. Feldman, P. J. Macguire, {\it Tridiagonal reproducing
%kernels and
%subnormality}, J. Operator Theory, {\bf 70} (2013), 477-494.

%\bibitem{AFM}
%G. Adams, N. Feldman, and P. McGuire, {Tridiagonal reproducing kernels and
%subnormality}, {\it J. Operator Theory}, {\bf 70} (2013), 477-494.

%\bibitem{AM-0}
%G. Adams and P. McGuire,
%Analytic tridiagonal reproducing kernels,
%{\it J. London Math. Soc.} {\bf 64} (2001), 722-738.

%\bibitem{AM-1} G.
%Adams and P. McGuire,
%A class of tridiagonal reproducing kernels,
%{\it Oper. Matrices} { 2} (2008), 233-247.



%\bibitem{AM} J. Agler and J. McCarthy, {\it Pick Interpolation and Hilbert Function Spaces}, Graduate Studies in Mathematics, 44. American Mathematical Society, Providence, RI, 2002, xx+308 pp.

\bibitem{AM}
J. Agler and J. McCarthy, Distinguished varieties, {\it Acta Math.} {\bf 194} (2005) 133–153.

%\bibitem{AY} J. Agler and N. Young, A model theory for $\Gamma$-contractions,
%{\it J. Operator Theory}, {\bf 49} (2003), 45-60.

%\bibitem{AP} E. Albrecht and M. Ptak,
%Invariant subspaces for doubly commuting contractions with rich Taylor spectrum,
%{\it J. Operator Theory}, {\bf 40} (1998), 373-384.

%\bibitem{Al} A. Aleman, {\it The multiplication operators on Hilbert spaces of
%analytic functions},
%Habilitationsschrift, Fernunuversitat Hagen, 1993.

%\bibitem{AMu} C. Ambrozie and V. M$\ddot{\mbox{u}}$ller,
%Invariant subspaces for polynomially bounded operators,
%{\it J. Funct. Anal.} {\bf 213} (2004), 321-345.

\bibitem{Ando} 
T. And$\hat{\mbox{o}}$, On a pair of commutative contractions, {\it Acta Sci. Math. (Szeged)}, {\bf 24} (1963), 88-90.

%\bibitem{ACJS} A. Anand, S. Chavan, Z. Jab{\l}o\'nski, and J. Stochel,
%Subnormality of the Cauchy Dual
%of a 2-isometry, {\it Preprint}.

%\bibitem{BJJS}
%P. Budzyński, Z. J. Jab{\l}o\'nski, I. B. Jung and J. Stochel,
%A subnormal weighted shift on a directed tree whose $n$th power has trivial %domain,
%{\it J. Math. Anal. Appl.} {\bf 435} (2016), 302-314.

%\bibitem{AHHK}
%W. Arveson, D. Hadwin, T. Hoover and E. Kymala,
%{Circular operators}, {\it Indiana Univ. Math. J.} {\bf 33} (1984), 583-595.

%\bibitem{AN} W. Arendt and N. Nikolski,
%Vector-valued holomorphic functions revisited,
%{\it Math. Z.} {\bf 234} (2000), 777-805.

%\bibitem{AL}  K. Attele and A. Lubin, Dilations and commutant lifting for jointly isometric operators$-$a geometric approach, {\it J. Funct. Anal.} {\bf 140} (1996), 300-311.

%\bibitem{BM-1} B. Bagchi and G. Misra, The homogeneous shifts, {\it J. Funct. Anal.} {\bf 204} (2003), 293-319.
%
%\bibitem{B}
%A. Bourhim, On the local spectral properties of weighted shift operators, {\it Studia Math.} {\bf 163} (2004), 41-69.
%
\bibitem{B}
F. Br$\ddot{\mbox{u}}$ckler, Tensor products of $C^*$-algebras, operator spaces and Hilbert $C^*$-modules, {\it Math. Commun.} {\bf 4} (1999), 257-268.
%
%\bibitem{B-D-P}  P. Budzy\'{n}ski, P. Dymek and M. Ptak,
%Analytic structure of weighted shifts on directed
%trees, {\em Math. Nachr.} {\bf 290} (2017), 1612-1629.

%\bibitem{BDPP} P. Budzy\'{n}ski,  P. Dymek, A. Planeta and M. Ptak, Weighted shifts on directed trees. Their multiplier algebras, reflexivity and decompositions, {\it Preprint}, (2017) arXiv:1702.00765 [math.FA]

%\bibitem{B0} H. Bercovici,
%The algebra of multiplication operators on Bergman spaces,
%{\it Arch. Math.}  {\bf 48} (1987), 165-174.
%
%\bibitem{B1} H. Bercovici,
%A factorization theorem with applications to invariant subspaces and the reflexivity of isometries, {\it Math. Res. Lett.} {\bf 1} (1994), 511-518.

%\bibitem{BFP} H. Bercovici, C. Foias, and C. Pearcy,
%Dual algebras with applications to invariant subspaces and dilation theory,
%CBMS Regional Conference Series in Mathematics, {\bf 56}, American Mathematical Society, Providence, RI, 1985.

%\bibitem{BC} S. Brown and B. Chevreau, Toute contraction \'{a} calcul fonctionnel isom\'{e}trique est r\'{e}flexive, {\it C. R. Acad. Sci. Paris S\'{e}r. I Math.} {\bf 307} (1988), 185-188.

\bibitem{CPT} S. Chavan, D. Pradhan and S. Trivedi, Multishifts on directed Cartesian products of rooted directed trees, {\it Dissertationes Math. (Rozprawy Mat.)} {\bf 527} (2017), 102 pp.

%\bibitem{CPT-1} S. Chavan, D. Pradhan and S. Trivedi, Classification of Drury-Arveson-type Hilbert modules associated with certain directed graphs, {\it Preprint}, (2017) arXiv:1709.02922 [math.FA]

%\bibitem{CT}
%S. Chavan and S. Trivedi, An analytic model for left-invertible weighted shifts on directed trees, {\it J. London Math. Soc.} {\bf 94} (2016), 253-279.

%\bibitem{CY}
%S. Chavan, D. Yakubovich, Spherical tuples of Hilbert space operators,
%{\it Indiana Univ. Math. J.} {\bf 64} (2015), 577-612.

%\bibitem{C-R} K. Clancey and D. Rogers, Cyclic vectors and seminormal
%operators, {\it Indiana Univ. Math. J.} {\bf 27} (1978), 689-696.

%\bibitem{CEP} B. Chevreau, G. Exner and C. Pearcy, On the structure of contraction operators III, {\it Michigan Math. J}, {\bf 36} (1989), 29-62.

%\bibitem{C-Z} M. Ch$\bar{\mbox{o}}$ and W. $\dot{\mbox{Z}}$elazko,
%{\it On the geometric radius of commuting $n$-tuples of operators}, Hokkaido Math. J. {\bf 21} (1992), 251-258.

%\bibitem{Conway}
%J. Conway, {\it A course in functional analysis}, Springer Verlag, New York,
%1985.

%\bibitem{Co} J. Conway, {\it The Theory of Subnormal Operators}, Math. Surveys Monographs,
%{\bf 36}, Amer. Math. Soc. Providence, RI 1991.

%\bibitem{Co1} J. Conway, {\it A Course in Operator Theory},  Graduate Studies in Mathematics, {\bf 21}, American Mathematical Society, Providence, RI, 2000. xvi+372 pp.

%\bibitem{CD} J. Conway and J. Dudziak, von Neumann operators are reflexive, {\it J. Reine Angew. Math.} {\bf 408} (1990), 34-56.


%\bibitem{C-D} M. Cowen, R. Douglas, Complex geometry and operator theory, {\it Acta Math.} {\bf 141}  (1978), 187-261.

\bibitem{CrD}
M. Crabb and A. Davie, Von Neumann's inequality for Hilbert space operators, {\it Bull. Lond. Math. Soc.} {\bf 7} (1975) 49-50.


\bibitem{Cu}
R. Curto, Applications of several complex variables to multiparameter spectral theory. Surveys of some recent results in operator theory, Vol. II, 25-90, {\it Pitman Res. Notes Math. Ser. {\bf 192}, Longman Sci. Tech.}, Harlow, 1988.

\bibitem{C-S} R. Curto and N. Salinas, Generalized Bergman kernels and the Cowen-Douglas theory, {\it Amer. J. Math.} {\bf 106} (1984), 447-488.

\bibitem{D}
S. Drury, Remarks on von {N}eumann's inequality, {\it Banach spaces, harmonic analysis, and probability theory (Storrs, Conn., 1980/1981)}, 14–32, 
Lecture Notes in Math., 995, Springer, Berlin, 1983.

%\bibitem{Di}  M. Didas,  Dual algebras generated by von Neumann $n$-tuples over strictly pseudoconvex sets, {\it Dissertationes Math. $($Rozprawy Mat.$)$} {\bf 425} (2004), 77 pp.
%
%\bibitem{Di-1} M. Didas,
%Spherical isometries are reflexive,
%{\it
%Integr. Eqn. Op. Th}, {\bf 52} (2005), 599-604.

%\bibitem{E} M. Engli$\check{\mbox{s}}$,
%Asymptotic behaviour of reproducing kernels, Berezin quantization and mean-value theorems, {\it Reproducing kernels and their applications}, (Newark, DE, 1997), 53–64,
%Int. Soc. Anal. Appl. Comput, {\bf 3}, Kluwer Acad. Publ., Dordrecht, 1999.

%\bibitem{Es0} J. Eschmeier,
%Multiplication operators on Bergman spaces are reflexive, Linear operators in function spaces (Timişoara, 1988), 165-184,
%{\it Oper. Theory Adv. Appl.} {\bf 43}, Birkh$\ddot{\mbox{a}}$user, Basel, 1990.

%\bibitem{Es1} J. Eschmeier, Invariant subspaces for spherical contractions, {\it Proc. London Math. Soc.} {\bf 75} (1997), 157-176.

%Invariant subspaces for commuting contractions, {\it J. Operator Theory} {\bf 45} (2001), 413-443.

%\bibitem{Es2} J. Eschmeier,
%$C_{00}$-representations of $H^{\infty}(G)$ with dominating Harte spectrum, Banach algebras '97 (Blaubeuren), 135-151, de Gruyter, Berlin, 1998.
%
%\bibitem{Es3} J. Eschmeier,
%Algebras of subnormal operators on the unit ball,
%{\it J. Operator Theory}, {\bf 42} (1999), 37-76.
%
%\bibitem{Es4}  J. Eschmeier,  Invariant subspaces for commuting contractions, {\it J. Operator Theory} {\bf 45} (2001), 413-443.

%\bibitem{Es}  J. Eschmeier, On the reflexivity of multivariable isometries, {\it Proc. Amer. Math. Soc.} {\bf 134} (2006), 1783-1789.

%\bibitem{G} 
%G. Geh\'{e}r, Asymptotic behaviour and cyclic properties of weighted shifts on directed trees, {\it J. Math. Anal. Appl.} {\bf 440} (2016), 14-32.
%
%\bibitem{Gel1}
%R. Gellar, {\it Shift operators in Banach space}, Dissertation, Columbia University,
%New York, N. Y., 1968.
%
%
%\bibitem{Gel2}
% R. Gellar, Operators commuting with a weighted shift, {\it Proc. Amer. Math. Soc.} {\bf 23} (1969), 538-545.
% 
%\bibitem{Gel3}
%R. Gellar, Cyclic vectors and parts of the spectrum of a weighted shift, {\it Trans. Amer. Math. Soc.} {\bf 146} (1969), 69-85.
%
%\bibitem{Ge} R. Gellar, {Circularly symmetric normal and subnormal operators}, {\it J. Analyse
%Math.} {\bf 32} (1977) 93-117.
%
%\bibitem{Gr}
%S. Grabiner, Weighted shifts and Banach algebras of power series, {\it Amer. J. Math.} {\bf 97} (1975), 16-42.

\bibitem{Thesis}
R. Gupta, The Carath\'{e}odory-Fej\'{e}r Interpolation Problems and the von-Neumann Inequality, Thesis, (2015), arXiv:1508.07199

\bibitem{Gu}
R. Gupta, An improvement on the bound for $C_2(n)$, {\it Acta Sci. Math. (Szeged)} {\bf 83} (2017), 263-269.

\bibitem{GR}
R. Gupta and S. Ray, On a Question of N. Th. Varopoulos
and the constant $C_2(n)$, {\it Ann. Inst. Fourier(Grenoble)}, to appear.

\bibitem{K}
G. Knese, The von {N}eumann inequality for {$3\times 3$} matrices, {\it Bull. Lond. Math. Soc.} {\bf 48} (2016), 53-57


\bibitem{Hz} M. Hartz, Von Neumann's inequality for commuting weighted shifts, {\it Indiana Univ. Math. J.} {\bf 66} (2017), 1065-1079.

%\bibitem{H}
%H. Helson, Invariant subspaces of the weighted shift. Hilbert space operators and operator algebras (Proc. Internat. Conf., Tihany, 1970), pp. 271–277. Colloq. Math. Soc. János Bolyai, 5, North-Holland, Amsterdam, 1972.
%
%\bibitem{Her}
%D. Herrero, Eigenvectors and cyclic vectors of bilateral weighted shifts. II. Simply invariant subspaces, {\it Integral Equations Operator Theory} {\bf 6} (1983), 515-524. 

\bibitem{Hol}
J. Holbrook, Schur norms and the multivariate von Neumann inequality, {\it Recent advances in operator theory and related topics (Szeged, 1999)}, (2001), 375-386.

%\bibitem{Hua}
%L. Hua, {\it Harmonic analysis of functions of several complex variables in the classical domains}. Translated from the Russian, which was a translation of the Chinese original, by Leo Ebner and Adam Korányi. With a foreword by M. I. Graev. Reprint of the 1963 edition. Translations of Mathematical Monographs, 6. American Mathematical Society, Providence, R.I., 1979. iv+186 pp. 

\bibitem{J}
Z. Jab{\l}o\'nski, Hyperexpansive operator valued unilateral weighted shifts, {\it Glasg. Math. J.} {\bf 46} (2004), 405-416.

%\bibitem{JJS}
%Z. Jab{\l}o\'nski, I. Jung and J. Stochel, Weighted shifts on directed
%trees, {\it  Mem. Amer. Math. Soc.} {\bf 216} (2012), no. 1017, viii+106.

%\bibitem{J} N. Jewell,
%Multiplication by the coordinate functions on the Hardy space of the unit sphere in $C^n$,
%{\it Duke Math. J.} {\bf 44} (1977), 839-851.

\bibitem{JL}
N. P. Jewell and A. R. Lubin, Commuting weighted shifts and analytic function theory in several variables, {\it J. Operator Theory}, {\bf 1} (1979), 207-223.


%\bibitem{KS} R. Khani and K. Seddighi,
%On the commutant of the direct sum of operators of multiplication by the independent variable, {\it J. Operator Theory}, {\bf 43} (2000),
%199-210.

%\bibitem{KP} M. Kosiek and M. Ptak, Reflexivity of N-tuples of contractions with rich joint left essential spectrum, {\it
%Integr. Eqn. Op. Th}, {\bf 13} (1990), 395-420.

%\bibitem{KO} M. Kosiek and A. Octavio,
%On common invariant subspaces for commuting contractions with rich spectrum,
%{\it Indiana Univ. Math. J.} {\bf 53} (2004), 823-844.

%\bibitem{KOP} M. Kosiek, A. Octavio and M. Ptak,
%On the reflexivity of pairs of contractions,
%{\it Proc. Amer. Math. Soc.} {\bf 123} (1995), 1229-1236.

%\bibitem{Kr} S. Krantz, {\it Function Theory of Several Complex Variables}, AMS Chelsea Publishing,
%Providence, RI, 2001.

%\bibitem{K}  T. Kriete,  Laplace transform asymptotics, Bergman kernels and composition operators, {\it Reproducing kernel spaces and applications}, 255-272, Oper. Theory Adv. Appl, {\bf 143}, Birkh$\ddot{\mbox{a}}$user, Basel, 2003.

%\bibitem{KM} A. Kor\'{a}nyi and G. Misra, Homogeneous operators on Hilbert spaces of holomorphic functions, {\it J. Funct. Anal.} {\bf 254} (2008),
%2419-2436.

%\bibitem{Ku} C. S. Kubrusly, {\it Spectral theory of operators on Hilbert
%spaces}, Birkhäuser, New York, 2012.

%\bibitem{K}
%S. Kumar, Spherically balanced Hilbert spaces of formal power series in several variables–II, {\it Complex Anal. Oper. Theory} {\bf 10} (2016), 505-526.

\bibitem{L} A. Lambert, Unitary equivalence and reducibility of invertibly weighted shifts, {\it Bull. Austral. Math. Soc.} {\bf 5} (1971), 157-173.

\bibitem{LT}
A. Lambert and T. Turner, The double commutant of invertibly weighted shifts, {\it Duke Math. J.} {\bf 39} (1972), 385-389.

%\bibitem{MS}  W. Majdak and J. Stochel, Weighted shifts on directed semi-trees: an application to creation operators on Segal-Bargmann spaces, {\it Complex Anal. Oper. Theory}, {\bf 10} (2016), 1427-1452.

%\bibitem{MMN}
%T. Miller, V. Miller and M. Neumann, Local spectral properties of weighted shifts, {\it J. Operator Theory} {\bf 51} (2004), 71-88.
%
%\bibitem{Ml} W. Mlak, {Note on circular operators. I, II, III, IV}, {\it Univ. Iagel. Acta Math.} {\bf 29} (1992) 145-152, 153-162, 163-170, 171-175.

%\bibitem{Mc} J. McCarthy, Reflexivity of subnormal operators, {\it Pacific J. Math.} {\bf 161} (1993), 359-370.

\bibitem{MS}
J. McCarthy, and O. Shalit, Unitary $N$-dilations for tuples of commuting matrices, {\it Proc. Amer. Math. Soc.} {\bf 141} (2013), 563-571.

%\bibitem{M-S} V. M$\ddot{\mbox{u}}$ller and A. Soltysiak, {\it Spectral radius formula for commuting Hilbert space operators},
%Studia Math. {\bf 103} (1992), 329-333.

%\bibitem{MV} V. M$\ddot{\mbox{u}}$ller and F.H. Vasilescu, {Standard
%models for some commuting multioperators}, \textit{Proc. Amer. Math. Soc.}
%\textbf{117} (1993), 979-989.

%\bibitem{N}
%N. Nikol$\acute{}$ski$\breve{\mbox{i}}$, {\it Treatise on the shift operator. Spectral function theory}. With an appendix by S. V. Hruščev [S. V. Khrushchëv] and V. V. Peller. Translated from the Russian by Jaak Peetre. Grundlehren der Mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences], 273. Springer-Verlag, Berlin, 1986. xii+491 pp.

%\bibitem{OT} R. Olin and J. Thomson, Algebras of subnormal operators, J. Funct. Anal. 37 (1980),
%271–301.

\bibitem{P} V. Paulsen, {\it Completely Bounded Maps and Operator Algebras}, Cambridge University Press, 2002.

%\bibitem{PR}  V. Paulsen and M. Raghupathi, {\it An Introduction to the Theory of Reproducing Kernel Hilbert Spaces}, Cambridge Studies in Advanced Mathematics, 152. Cambridge University Press, Cambridge, 2016.

%\bibitem{Pi}  G. Pisier, {\it  Similarity Problems and Completely Bounded Maps}, Second, expanded edition; Includes the solution to “The Halmos problem”, Lecture Notes in Mathematics, vol. 1618, Springer-Verlag, Berlin, 2001.

%\bibitem{Pr}  B. Prunaru, von Neumann operators are reflexive, {\it
%Integr. Eqn. Op. Th}, {\bf 14} (1991), 609-612.
%\bibitem{R} W. Ridge, {\it Approximate point spectrum of a weighted
%shift}, Tran. Amer. Math. Soc. {\bf 147}, (1970), 349-356.

%\bibitem{Pi}  G. Pisier, {\it Introduction to operator space theory}, London Mathematical Society Lecture Note Series, 294. Cambridge University Press, Cambridge, 2003. viii+478 pp.

\bibitem{Pt}  M. Ptak, On the reflexivity of multigenerator algebras, {\it Dissertationes Math. $($Rozprawy Mat.$)$} {\bf 378} (1998), 61 pp.

%\bibitem{RR}  H. Radjavi and P. Rosenthal, On invariant subspaces and reflexive algebras, {\it Amer. J. Math.} {\bf 91} (1969), 683-692.

%\bibitem{OT} R. Olin and J. Thomson,
%Algebras of subnormal operators,
%{\it J. Funct. Anal.} {\bf 37} (1980), 271-301.
%
%\bibitem{RR} H. Radjavi and P. Rosenthal,  {\it Invariant subspaces}, Second edition. Dover Publications, Inc., Mineola, NY, 2003. xii+248 pp.
%
%\bibitem{R} R. Range, {\it Holomorphic Functions and Integral Representations in Several Complex
%Variables}, Springer-Verlag, New York, 1986.

%\bibitem{Re} O. R\'{e}jasse, Factorization and reflexivity results for polynomially bounded operators, {\it J. Operator Theory} {\bf 60} (2008), 219-238.

%\bibitem{Ri} S. Richter, Invariant subspaces of the Dirichlet shift, {\it J. Reine Angew. Math.} {\bf 386} (1988), 205-220.
%
%\bibitem{Sa0}  D. Sarason, Invariant subspaces and unstarred operator algebras, {\it Pacific J. Math.} {\bf 17} (1966), 511-517.
%
%\bibitem{Sa} D. Sarason,
%{Generalized interpolation in $H^{\infty}$},
%{\it Trans. Amer. Math. Soc.} {\bf 127} (1967), 179-203.

\bibitem{Sarkar}
B. Das, J. Sarkar and S. Sarkar, Factorizations of contractions,  {\it Adv. Math.} {\bf 322} (2017), 186-200.

%
%\bibitem{Se}  K. Seddighi, Reproducing kernel Hilbert spaces, {\it Iranian J. Sci. Tech.} {\bf 17} (1993), 171-177.

\bibitem{S} A. Shields, {\it Weighted shift operators and analytic function
theory, in Topics in Operator Theory}, Math. Surveys
Monographs, vol. 13, Amer. math. Soc., Providence, RI 1974, 49-128.

%\bibitem{Sh-Wa} A. Shields and L. Wallen,
%The commutants of certain Hilbert space operators,
%{\it Indiana Univ. Math. J.} {\bf 20} (1971), 777-788.

%\bibitem{Si}  B. Simon, {\it Real analysis. A Comprehensive Course in Analysis}, Part 1. American Mathematical Society, Providence, RI, 2015.

%\bibitem{Shimorin}
%S. Shimorin, Wold-type decompositions and wandering subspaces for operators
%close to isometries, {\it J. Reine Angew. Math.} {\bf 531} (2001), 147-189.
\bibitem{Sz}
B. Sz.-Nagy, Sur les contractions de l’espace de Hilbert, {\it Acta Sci. Math. (Szeged)} {\bf 15} (1953) 87-92.

\bibitem{SF} 
B. Sz$\ddot{\mbox{o}}$kefalvi-Nagy, C. Foias,
H. Bercovici and L. K\'erchy, {\it Harmonic analysis of
operators on Hilbert space}, Springer, New York, 2010.

%\bibitem{SS}
%J. Stochel and J. B. Stochel, On the $k$th root of a Stieltjes moment sequence, {\it J. Math. Anal. Appl.} {\bf 396} (2012), 786-800.

%\bibitem{SV} V. N. Singh, Reproducing kernels and operators with a cyclic vector. I. {\it Pacific J. Math.} {\bf 52} (1974), 567-584.

\bibitem{V1}
N. Varopoulos, On an inequality of von Neumann and an application of the metric theory of tensor products to operators theory, {\it J. Functional Analysis} {\bf 16} (1974), 83-100.

\bibitem{V2}
N. Varopoulos, On a commuting family of contractions on a Hilbert space, {\it Rev. Roumaine Math. Pures Appl.} {\bf 21} (1976), 1283-1285.

\bibitem{von}
J. von Neumann, Eine Spektraltheorie fur allgemeine Operatoren eines unitdren Raumes, {\it Math. Nachr.} {\bf 4} (1951) 258-281.

%\bibitem{Y-2} D. Yakubovich, Dual piecewise analytic bundle shift models of linear operators, {\it J. Funct. Anal.}
%{\bf 136}  (1996), 294-330.

%\bibitem{Y} T. Yoshino,
%Subnormal operator with a cyclic vector,
%{\it T$\hat{\mbox{o}}$hoku Math. J.} {\bf 21} (1969), 47-55.


\end{thebibliography}

\documentclass[a4paper,oneside,10pt, reqno]{amsproc}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{mathrsfs}
\usepackage{epsfig}
\usepackage{graphicx}
\usepackage{epstopdf}
\usepackage{color}
\usepackage{ifthen}
\usepackage{float}

\makeatletter
\@namedef{subjclassname@2010}{%
\textup{2010} Mathematics Subject Classification}
\makeatother

\newtheorem{theorem}{Theorem}[section]
\newtheorem{proposition}[theorem]{Proposition}%[section]
\newtheorem{corollary}[theorem]{Corollary}%[section]
\newtheorem{lemma}[theorem]{Lemma}%[section]

\theoremstyle{remark}
\theoremstyle{definition}
\newtheorem{example}[theorem]{Example}

\newcommand{\bq}{\begin{equation}}
\newcommand{\eq}{\end{equation}}
\newcommand{\beqn}{\begin{eqnarray*}}
\newcommand{\eeqn}{\end{eqnarray*}}
\newcommand{\beq}{\begin{eqnarray}}
\newcommand{\eeq}{\end{eqnarray}}

\newcommand{\rar}{\rightarrow}

\newcommand{\bc}{\begin{centre}}
\newcommand{\ec}{\end{centre}}

\newcommand{\ba}{\begin{array}}
\newcommand{\ea}{\end{array}}

\newcommand{\inp}[2]{\langle{#1},\,{#2} \rangle}

\renewcommand{\Delta}{{\nabla}}

\newcommand*{\Ge}{\geqslant}

\newcommand*{\Le}{\leqslant}

\begin{document}
\title[Von Neumann's inequality for operator-valued multishifts]
{Von Neumann's inequality for commuting \\operator-valued multishifts}
\author[R. Gupta]{Rajeev Gupta}
\author[S. Kumar]{Surjit Kumar}
\author[S. Trivedi]{Shailesh Trivedi}
\address{Department of Mathematics and Statistics\\
Indian Institute of Technology  Kanpur, India}
   \email{rajeevg@iitk.ac.in}
\address{Department of Mathematics \\ Indian Institute of Science Bangalore, India}
\email{surjitkumar@iisc.ac.in}
\address{Department of Mathematics and Statistics\\
Indian Institute of Technology  Kanpur, India}
   \email{shailtr@iitk.ac.in}

\thanks{The work of first and second author was supported by Inspire Faculty Fellowship (Ref. No. DST/INSPIRE/04/2017/002367, DST/INSPIRE/04/2016/001008) while that of third author was supported by the National Post-doctoral Fellowship (Ref. No. PDF/2016/001681), SERB}

%   \thanks{The research of the
%third and fourth authors was supported by the NCN
%(National Science Center), decision No.
%DEC-2013/11/B/ST1/03613.}
   \subjclass[2010]{Primary 47B37, Secondary 47A13}
\keywords{operator-valued multishift, von Neumann's inequality, tensor product}

\date{}

\begin{abstract}
Recently, Hartz proved that every commuting contractive classical multishift with non-zero weights satisfies the matrix-version of von Neumann's inequality. We show that this result does not extend to the class of commuting operator-valued multishifts with invertible operator weights. In particular, we show that if $A$ and $B$ are commuting contractive $d$-tuples of operators such that $B$ satisfies the matrix-version of von Neumann's inequality and $(1, \ldots, 1)$ is in the algebraic spectrum of $B$, then the tensor product $A \otimes B$ satisfies the von Neumann's inequality if and only if $A$ satisfies the von Neumann's inequality. We also exhibit several families of operator-valued multishifts for which the von Neumann's inequality always holds. 
\end{abstract}

\maketitle

\section{Introduction}

The celebrated von Neumann's inequality \cite{von} says that if $T$ is a contraction on a Hilbert space $\mathcal H$, then $\|p(T)\| \Le \sup_{|z|<1} |p(z)|$ for every polynomial $p$. Generalizing this result, Sz.-Nagy \cite{Sz} proved that every contraction has a unitary dilation. Later Ando \cite{Ando} (see also \cite{SF}) extended this result and showed that every pair of commuting contractions dilates to a pair of commuting unitaries, and hence, every pair of commuting contractions satisfies the von Neumann's inequality. Thus it is natural to ask whether the von Neumann's inequality holds for a $d$-tuple of commuting contractions, $d \Ge 3$. Surprisingly, it fails for $d \Ge 3$. In fact, Varopoulos, in \cite{V1}, showed that there exists a big enough $d$ for which the von Neumann's inequality fails for a $d$-tuple of commuting contractions. In the addendum of the same paper, he together with Kaijser and independently Crabb and Davie \cite{CrD} gave examples of three commuting contractions which do not satisfy the von Neumann's inequality. Since then it has been one of the peculiar topics in operator theory.  In \cite[Question 36]{S}, Shields asked whether a $d$-tuple of commuting contractive weighted shifts (in other words, contractive classical multishift) satisfies the von Neumann's inequality. This question was attributed to Lubin and was explicitly mentioned in \cite{JL}. Recently, Hartz \cite{Hz} answered this question affirmatively and proved the following result:
\begin{theorem}\cite[Theorem 1.1]{Hz}\label{Hartz}
Let $T = (T_1, \ldots, T_d)$ be a contractive classical multishift with non-zero weights. Then $T$ dilates to a $d$-tuple of commuting unitaries.
\end{theorem}
In view of this, it is natural to ask whether the above result extends to the class of commuting operator-valued multishifts with invertible operator weights. The purpose of this note is to study the von Neumann's inequality for commuting operator-valued multishifts. A key tool in this study is the following  characterization for the tensor product of two $d$-tuples of commuting contractions to satisfy the von Neumann's inequality.

\begin{theorem}\label{intro-thm}
Let $d$ be a positive integer and let $A = (A_1, \ldots, A_d)$, $B = (B_1, \ldots, B_d)$ be two commuting $d$-tuples of contractions on the Hilbert spaces $\mathcal H$ and $\mathcal K$ respectively. Suppose that $B$ satisfies the matrix-version of von Neumann's inequality and $(1, \ldots, 1)$ belongs to the algebraic spectrum $\sigma(B)$ of $B$. Then $A \otimes B = (A_1 \otimes B_1, \ldots, A_d \otimes B_d)$ satisfies the von Neumann's inequality if and only if $A$ satisfies the von Neumann's inequality. 
\end{theorem}

Using this characterization, we prove that if $A = (A_1, \ldots, A_d)$ is a $d$-tuple of commuting contractions on a Hilbert space $H$ and $T = (T_1, \ldots, T_d)$ is a commuting operator-valued multishift on $\ell^2_H(\mathbb N^d)$ with operator weights given by $A^{(j)}_\alpha = A_j$ for all $\alpha \in \mathbb N^d$ and $j = 1, \ldots, d$, then $T$ satisfies the von Neumann's inequality if and only if $A$ satisfies the von Neumann's inequality. This readily yields a family of operator-valued multishifts with invertible operator weights which do not satisfy the von Neumann's inequality. This is in contrast with Theorem \ref{Hartz}. We conclude this paper with a concrete example of a commuting operator-valued multishift with invertible operator weights which does not satisfy the von Neumann's inequality. This example is motivated by the one which Kaijser and Varopoulos \cite{V1} gave to disprove the von Neumann's inequality for $3$-tuple of commuting contractions. It is worth mentioning that a commuting $d$-tuple of contractions dilates to a commuting $d$-tuple of unitaries if and only if it satisfies the matrix version of von Neumann's inequality \cite[Corollary 7.7]{P}. We refer the reader to \cite{P, AM, SF, MS, Sarkar} for recent developments related to the von Neumann's inequality.

We set below the notations used in posterior sections.
For a set $X$ and a positive integer $d$,  $X^d$ stands for the $d$-fold Cartesian product of $X$.
The symbols ${\mathbb N}, \ \mathbb Z,\ \mathbb R$ and $\mathbb C$ stand for the set of nonnegative
integers, set of integers, the field of real numbers and the field of complex numbers, respectively.
For $\alpha =
(\alpha_1, \ldots, \alpha_d) \in {\mathbb{N}}^d,$
we set $|\alpha|:=\sum_{j=1}^d
\alpha_j$.
For $w=(w_1, \ldots, w_d) \in \mathbb C^d$ and $\alpha=(\alpha_1, \ldots, \alpha_d) \in \mathbb N^d$, the complex conjugate $\overline{w} \in \mathbb C^d$ of $w$ is given by $(\overline{w}_1, \ldots, \overline{w}_d),$ while $w^\alpha$ denotes the complex number $\prod_{j=1}^d w^{\alpha_j}_j$. The symbol $\mathbb D^d$ is reserved for the open unit polydisc in $\mathbb C^d$ centered at the origin whereas $\overline{\mathbb D}^d$ stands for the closed unit polydisc in $\mathbb C^d$. 
Let $\mathcal H$ be a complex Hilbert space.
If $F$ is a subset of $\mathcal H$, the closed linear span of $F$ is denoted by $\bigvee \{x : x \in F\}$.  For a positive integer $m,$ the orthogonal direct
sum of $m$ copies of $\mathcal H$ is denoted by $\mathcal H^{(m)}.$ 
Let $\mathcal{B}({\mathcal H})$ denote the unital Banach algebra of
bounded linear operators on $\mathcal H.$ The multiplicative identity $I$ of 
$\mathcal{B}(\mathcal H)$ is sometimes denoted by $I_{\mathcal H}$.
The norm on  $\mathcal H$ is denoted by $\|\cdot\|_{\mathcal H}$ 
and whenever there is no confusion likely, 
we remove the subscript $\mathcal H$ from $\|\cdot\|_{\mathcal H}$. 
If $T \in \mathcal B(\mathcal H)$, then $T^*$ denotes the Hilbert space adjoint of $T$. An operator $T \in \mathcal B(\mathcal H)$ is said to be a {\it contraction} if $\|T\| \Le 1$. 
By a  {\it commuting $d$-tuple $T=(T_1, \ldots, T_d)$ in $\mathcal B(\mathcal H)$}, 
we mean a collection of commuting operators $T_1, \ldots, T_d$ in $\mathcal B(\mathcal H)$ and $T$ is said to be {\it contractive} if $T_j$ is a contraction for each $j=1, \ldots, d.$  
For $\alpha = (\alpha_1, \ldots, \alpha_d) \in \mathbb N^d$, we understand $T^{\alpha}$ as the operator $T^{\alpha_1}_1\cdots T^{\alpha_d}_d$, where we adhere to the convention that $A^0 = I_\mathcal H$ for $A \in  \mathcal B(\mathcal H)$. 
Let $T = (T_1, \ldots, T_d)$ be a commuting $d$-tuple in $\mathcal B(\mathcal H)$ and $\mathfrak B$ be the unital Banach algebra generated by $T_1, \ldots, T_d$. Then the algebraic spectrum $\sigma(T)$ of $T$ is given by
\beqn
\sigma(T) = \mathbb C^d \setminus \Big\{\lambda \in \mathbb C^d : \mbox{ there exist } S_1, \ldots, S_d \in \mathfrak B \mbox{ such that } \sum_{j=1}^d (T_j - \lambda_j)S_j = I_\mathcal H\Big\}.
\eeqn
The reader is referred to \cite{Cu} for a detailed account on various notions of spectra of a commuting tuple of operators.   
A $d$-tuple $T = (T_1, \ldots, T_d)$ of commuting contractions on $\mathcal H$ is said to satisfy the matrix-version of von Neumann's inequality if for every positive integer $m$,
\beqn
\|(p_{i, j}(T))_{_{1 \Le i, j \Le m}}\|_{\mathcal B(\mathcal H^{(m)})} \Le  \sup_{z \in \mathbb D^d}\|(p_{i, j}(z))_{_{1 \Le i, j \Le m}}\|_{\mathcal B(\mathbb C^m)}, \quad p_{i, j} \in \mathbb C[z_1, \ldots, z_d],
\eeqn
where $\mathbb C[z_1, \ldots, z_d]$ denotes the ring of polynomials over $\mathbb C$ in $d$ complex variables $z_1, \ldots, z_d$. 

\section{Von Neumann's inequality for tensor product of tuples}
 
In this section, we prove Theorem \ref{intro-thm}. We are grateful to the anonymous referee for his constructive comments which significantly improved the earlier version of this theorem. 
 
%\begin{lemma}\label{spectral-radius}
%Let $d$ be a positive integer and $A = (A_1, \ldots, A_d)$, $B = (B_1, \ldots, B_d)$ be two $d$-tuples of commuting contractions on the Hilbert spaces $\mathcal H$ and $\mathcal K$ respectively. If $(1, \ldots, 1)$ is in the algebraic spectrum $\sigma(B)$ of $B$, then for a polynomial $p \in \mathbb C[z_1, \ldots, z_d]$, 
%\end{lemma}
%
%\begin{proof}
%Let $p \in \mathbb C[z_1, \ldots, z_d]$ be a polynomial. Then by the spectral mapping theorem \cite{Cu}, we get $p(\sigma(B)) = \sigma(p(B))$. Hence
%\beqn
%r(p(B)) = \sup_{\lambda \in \sigma(p(B))} |\lambda| = \sup_{\lambda \in p(\sigma(B))} |\lambda| = \sup_{z \in \sigma(B)} |p(z)|.
%\eeqn
%Since $\mathbb T^d \subseteq \sigma(B) \subseteq \overline{\mathbb D}^d$, it follows that $\sup_{z \in \sigma(B)} |p(z)| = \sup_{z\in {\mathbb D}^d} |p(z)|$. This gives the desired conclusion.
%\end{proof}
%
%We are now ready to prove the Theorem \ref{intro-thm}.

\begin{proof}[Proof of Theorem \ref{intro-thm}]
Since $(1, \ldots, 1) \in \sigma(B)$, it follows from \cite[Proposition 1.2]{Cu} that there exists a multiplicative linear functional $\chi$ on the unital Banach algebra $\mathfrak B$ generated by $B_1, \ldots, B_d$ such that $\chi(B_j) = 1$ for all $j = 1, \ldots, d$.
Let $k\in\mathbb N$ and 
\beqn
p(z)=\sum_{\underset{|\alpha| \Le k}{\alpha \in \mathbb N^d}}a_{\alpha}z^\alpha, \quad a_\alpha\in\mathbb C,\ z \in \mathbb C^d,
\eeqn
be a polynomial. 
%Consider the matrix valued polynomial $p_A(z)$ given by 
%\beqn
%p_{_A}(z)=\sum_{\underset{|\alpha| \Le k}{\alpha \in \mathbb N^d}}a_{\alpha}A^\alpha z^\alpha, \ z \in \mathbb C^d.
%\eeqn
%Since $\mathcal B(\mathcal H) \otimes \mathcal B(\mathcal K)$ is isometrically embedded in $\mathcal B(\mathcal H \otimes \mathcal K)$, it follows that 
%\beq\label{p-tensor}
%p_{_A}(B) = p(A \otimes B). \textcolor{red}{??}
%\eeq
By Hahn-Banach theorem there exists a contractive linear functional $\phi$ on $\mathcal B(\mathcal H)$ such that $\phi(p(A)) = \|p(A)\|$. Then $\phi \otimes \chi$ is a contractive linear functional on $\mathcal B(\mathcal H) \otimes \mathfrak B$ (see \cite{B}). Hence, we get
\beqn
\|p(A \otimes B)\| \Ge |(\phi \otimes \chi) p(A \otimes B)| = \Big|\sum_{\underset{|\alpha| \Le k}{\alpha \in \mathbb N^d}}a_{\alpha}\phi(A^\alpha) \chi(B^\alpha) \Big| = |\phi(p(A))| = \|p(A)\|.
\eeqn
This establishes that $A$ satisfies the von Neumann's inequality if $A \otimes B$ satisfies the von Neumann's inequality. 

Conversely, assume that $A$ satisfies the von Neumann's inequality. Let $k\in\mathbb N$ and 
\beqn
p(z)=\sum_{\underset{|\alpha| \Le k}{\alpha \in \mathbb N^d}}a_{\alpha}z^\alpha, \quad a_\alpha\in\mathbb C,\ z \in \mathbb C^d,
\eeqn
be a polynomial.  Consider the polynomial $p(z,w)$ given by 
\beqn
p(z,w)=\sum_{\underset{|\alpha| \Le k}{\alpha \in \mathbb N^d}} a_{\alpha} z^\alpha w^\alpha,  \quad z, w \in \mathbb C^d.
\eeqn
Since $A$ satisfies the von Neumann's inequality, for each fixed $w \in \mathbb C^d$, we get
\beq\label{6}
\|p(A,w)\|=\Big\|\sum_{\underset{|\alpha| \Le k}{\alpha \in \mathbb N^d}}a_\alpha A^\alpha w^\alpha \Big\| \Le \sup_{z\in {\mathbb D}^d}|p(z,w)|.
\eeq
Let $\{e_\lambda : \lambda \in \Lambda\}$ be an orthonormal basis of $\mathcal H$. For a finite subset $F$ of $\Lambda$, let $P_F$ denote the orthogonal projection of $\mathcal H$ onto the subspace $\bigvee\{e_\lambda : \lambda \in F\}$ of $\mathcal H$. Define the (matrix-valued) polynomial
\beqn
p_{_F}(A, w) = \sum_{\underset{|\alpha| \Le k}{\alpha \in \mathbb N^d}}a_\alpha\, P_FA^\alpha  P_F \, w^\alpha, \quad w \in \mathbb C^d.
\eeqn 
Note that for each $w \in \mathbb C^d$, the net $p_{_F}(A, w)$ converges to $p(A, w)$ in the strong operator topology and $\|p_{_F}(A, w)\| \Le \|p(A, w)\|$. Since $B = (B_1, \ldots, B_d)$ satisfies the matrix version of von Neumann's inequality, it follows that
\beq\label{star}
\|p_{_F}(A, B)\| \Le \sup_{w \in \mathbb D^d} \|p_{_F}(A, w)\| \Le \sup_{w \in \mathbb D^d} \|p(A,w)\|.
\eeq
Also observe that the net $p_{_F}(A, B)$ converges to $p(A, B)$ in strong operator topology. Since $\mathcal B(\mathcal H) \otimes \mathcal B(\mathcal K)$ is isometrically embedded in $\mathcal B(\mathcal H \otimes \mathcal K)$, we identify $p(A, B)$ with $p(A \otimes B)$. Therefore, we get 
\beqn
\|p(A\otimes B)\| \overset{\eqref{star}}\Le \sup_{w \in \mathbb D^d} \|p(A,w)\|\overset{\eqref{6}}\Le \sup_{w\in {\mathbb D}^d} \sup_{z\in\mathbb D^d}|p(z,w)|= \sup_{z\in\mathbb D^d}|p(z)|.
\eeqn
%
%Observe that $p(A, B)=p(A \otimes B)$ and 
%since $B = (B_1, \ldots, B_d)$ satisfies matrix version of the von Neumann's inequality, it follows from \eqref{6} that 
%\beqn
%\|p(A \otimes B)\|_{\mathcal B(\mathbb C^n \otimes \mathcal H)}&=&\|p_{B}(A)\|_{\mathcal B(\mathbb C^n \otimes \mathcal H)} \Le \sup_{z\in\mathbb D^d}\|p_{z}(A)\|_{\mathcal B(\mathbb C^n)}\\
%&\Le& \sup_{z\in\mathbb D^d}\sup_{w\in {\mathbb D}^d}|p_{z}(w)|= \sup_{z\in\mathbb D^d}|p(z)|.
%\eeqn
This completes the proof of the theorem.
\end{proof}

\section{Operator-valued multishift and the von Neumann's inequality}

This section is devoted to the study of the von Neumann's inequality for commuting operator-valued multishifts. Before exhibiting a family of commuting contractive operator-valued multishifts with invertible operator weights which do not satisfy the von Neumann's inequality, we briefly recall the notion of operator-valued multishift. The notion of {\it operator-valued unilateral weighted shift} was introduced by Lambert in \cite{L} and was studied considerably thereafter (see \cite{LT, J} for related study). We refer to its several variable generalization as the {\it operator-valued multishift}. It seems that the notion of operator-valued multishift was not formally introduced and systematically studied earlier but it appeared at several places in the literature, see for instance \cite{C-S}, \cite{Pt}, \cite{CPT}. We now proceed towards the formal definition of operator-valued multishift.

Let $d$ be a positive integer and $\{H_\alpha : \alpha \in \mathbb N^d\}$ be a multisequence of complex Hilbert spaces. Let ${\mathcal H}=\oplus_{\alpha \in \mathbb N^d} H_\alpha$ be the orthogonal direct sum of $H_\alpha$, $\alpha \in \mathbb N^d$. Then $\mathcal H$ is a Hilbert space with respect to the following inner product:
\beqn \inp{x}{y}_{_{\mathcal H}}= \sum_{\alpha \in \mathbb N^d} \inp{x_{\alpha}}{y_{\alpha}}_{_{H_\alpha}},
\quad x=\oplus_{\alpha \in \mathbb N^d}x_{\alpha},\
y=\oplus_{\alpha \in \mathbb N^d} y_{\alpha} \in {\mathcal H}.\eeqn
If $H_\alpha = H$ for all $\alpha \in \mathbb N^d$, then we denote ${\mathcal H}=\oplus_{\alpha \in \mathbb N^d} H$ by $\ell^2_{H}(\mathbb N^d)$.
Let $\{{A^{(j)}_{\alpha}} : \alpha \in  \mathbb N^d,\ j = 1, \ldots, d \}$ be a multisequence of bounded linear operators $A^{(j)}_\alpha : H_\alpha \rar H_{\alpha+\varepsilon_j},$ where $\varepsilon_j$ is the $d$-tuple in $\mathbb N^d$ with $1$ in the $j^{\mbox{\tiny{th}}}$ place and zeros
elsewhere. 
An {\it operator-valued multishift} $T$ on $\mathcal H = \oplus_{\alpha \in \mathbb N^d} H_\alpha$ with operator weights $\{{A^{(j)}_{\alpha}} : \alpha \in  \mathbb N^d,\ j = 1, \ldots, d \}$ is a $d$-tuple of operators $T_1, \ldots, T_d$ in $\mathcal H$ defined by
\beqn 
\mathcal D(T_j) &:=& \Big \{ x=\oplus_{\alpha \in \mathbb N^d }x_{\alpha} \in {\mathcal H} :
\sum_{\alpha \in \mathbb N^d} \|A^{(j)}_{\alpha}x_{\alpha} \|^2 < \infty\Big\},\\
 T_j(\oplus_{\alpha \in \mathbb N^d}x_{\alpha})&:=& \oplus_{\alpha \in \mathbb N^d}
A^{(j)}_{\alpha-\varepsilon_j}x_{\alpha-\varepsilon_j}, \quad x=\oplus_{\alpha \in \mathbb N^d}x_{\alpha}
\in \mathcal D(T_j), \ j=1, \ldots, d.\eeqn 
For each $\alpha \in \mathbb N^d$ and $j=1,\ldots, d,$ if $\alpha_j=0$, then we
interpret $A^{(j)}_{\alpha-\varepsilon_j}$ to be a zero operator and $x_{\alpha-\varepsilon_j}$ as a zero vector.
Note that each $T_j$, $j=1, \ldots, d$, is a densely defined linear operator in $\mathcal H$.

The following proposition studies the basic properties of an operator-valued multishift. Its proof is straightforward and is left for the interested readers.  

\begin{proposition}
Let $d$ be a positive integer and $T = (T_1, \ldots, T_d)$ be an operator-valued multishift on $\mathcal H = \oplus_{\alpha \in \mathbb N^d} H_\alpha$ with operator weights $\{{A^{(j)}_{\alpha}} : \alpha \in  \mathbb N^d,\ j = 1, \ldots, d \}$. Then the following statements hold:
\begin{enumerate}
\item[(i)] For $j=1,\ldots,d$, $T_j$ is bounded if and only if
\beq \label{tj-bdd}
\sup_{\alpha \in \mathbb N^d} \|A^{(j)}_\alpha\| < \infty.
\eeq
%Further, if $T_j$ is bounded, then $\|T_j\| = \displaystyle\sup_{\alpha \in \mathbb N^d} \|A^{(j)}_\alpha\|$.
\item[(ii)] For $i,j=1, \ldots, d$, $T_i$ commutes with $T_j$ if and only if
\beq \label{commuting}
 A^{(i)}_{\alpha+\varepsilon_j} A^{(j)}_{\alpha}= A^{(j)}_{\alpha+\varepsilon_i} A^{(i)}_{\alpha} \ \mbox{for all } \alpha \in \mathbb N^d.
\eeq
\end{enumerate}
\end{proposition}

Let $T = (T_1, \ldots, T_d)$ be an operator-valued multishift on $\mathcal H = \oplus_{\alpha \in \mathbb N^d} H_\alpha$ with operator weights
$\{A^{(j)}_{\alpha}: \alpha \in \mathbb N^d,\  j=1, \ldots, d\}$. We refer to $T$ as {\it commuting operator-valued multishift} if the operator weights satisfy \eqref{tj-bdd} and \eqref{commuting}. Let us see how the class of classical multishifts is contained in that of operator-valued multishifts.

Let $\{w^{(j)}_\alpha : \alpha \in \mathbb N^d,\ j=1, \ldots,d\}$ be a multisequence of non-zero complex numbers such that $\sup_{\alpha \in \mathbb N^d} |w^{(j)}_\alpha| < \infty$ and $w^{(j)}_{\alpha+\varepsilon_i} w^{(i)}_\alpha = w^{(i)}_{\alpha+\varepsilon_j} w^{(j)}_\alpha$ for all $\alpha \in \mathbb N^d$, $i,j=1,\ldots,d$. Let $\mathcal H = \ell^2_{\mathbb C}(\mathbb N^d)$. Set $A^{(j)}_\alpha := w^{(j)}_\alpha I_{\mathbb C}$ for all $\alpha \in \mathbb N^d$ and $j=1,\ldots,d$. Then the commuting operator-valued multishift $T=(T_1, \ldots, T_d)$ with operator weights $\{A^{(j)}_{\alpha}: \alpha \in \mathbb N^d,\  j=1, \ldots, d\}$ is commonly known as {\it classical multishift} which was introduced in \cite{JL}. Note that $\ell^2_{\mathbb C}(\mathbb N^d)$ is nothing but $\ell^2(\mathbb N^d)$ and hence, in future, we shall write $\ell^2(\mathbb N^d)$ in place of $\ell^2_{\mathbb C}(\mathbb N^d)$.

We produce several families of commuting operator-valued multishifts for which the von Neumann's inequality always holds. We begin with the following lemma which generalizes \cite[Corollary 3.2]{L}.

\begin{lemma} \label{unit-wt-eq}
Let $d$ be a positive integer and $H$ be a complex Hilbert space. Let $T = (T_1, \ldots, T_d) \mbox { and } \tilde{T} = (\tilde{T_1}, \ldots, \tilde{T_d})$ be two commuting operator-valued multishifts on $\ell^2_{H}(\mathbb N^d)$ with respective unitary operator weights
$\{A^{(j)}_{\alpha}: \alpha \in \mathbb N^d, \  j=1, \ldots, d\}$ and $\{\tilde{A}^{(j)}_{\alpha}: \alpha \in \mathbb N^d, \  j=1, \ldots, d\}$. Then $T$ and $\tilde{T}$ are unitarily equivalent.
\end{lemma}

\begin{proof}
Let $U_0 = I_H$ and $U_{\alpha + \epsilon_j} = \tilde{A}^{(j)}_{\alpha} U_\alpha A^{(j)*}_{\alpha}$ for all $\alpha \in \mathbb N^d$ and $j = 1, \ldots, d$. We show that $U_\alpha$ is well-defined for each $\alpha \in \mathbb N^d$. To this end, first note that if $\alpha \in \mathbb N^d$ is such that $|\alpha| \Le 1$, then $U_\alpha$ is well defined. Now assume that   $|\alpha| \Ge 2.$ Let $\alpha = \beta + \epsilon_j = \gamma + \epsilon_k = \delta + \epsilon_j + \epsilon_k$ for some $\beta, \gamma, \delta \in \mathbb N^d$ and $j,k \in \{1, \ldots, d\}$. Using \eqref{commuting}, we get
\beqn
U_{\beta+\epsilon_j} &=&  \tilde{A}^{(j)}_{\beta} U_\beta A^{(j)*}_{\beta} = \tilde{A}^{(j)}_{\delta+\epsilon_k} U_{\delta+\epsilon_k} A^{(j)*}_{\delta+\epsilon_k} = \tilde{A}^{(j)}_{\delta+\epsilon_k} \tilde{A}^{(k)}_{\delta} U_\delta A^{(k)*}_{\delta} A^{(j)*}_{\delta+\epsilon_k}\\
&=& \tilde{A}^{(k)}_{\delta+\epsilon_j} \tilde{A}^{(j)}_{\delta} U_\delta A^{(j)*}_{\delta} A^{(k)*}_{\delta+\epsilon_j} = \tilde{A}^{(k)}_{\delta+\epsilon_j} U_{\delta+\epsilon_j} A^{(k)*}_{\delta+\epsilon_j} = \tilde{A}^{(k)}_{\gamma} U_\gamma A^{(k)*}_{\gamma}\\
&=& U_{\gamma+\epsilon_k}.
\eeqn
Since $\tilde{A}^{(j)}_{\alpha}$ and $A^{(j)}_{\alpha}$ are unitary operators,  it is clear that each $U_\alpha$ is a unitary operator on $H$. Set $U := \oplus_{\alpha \in \mathbb N^d} U_\alpha$. It is a routine verification to show that $U$ is a unitary on $\ell^2_{H}(\mathbb N^d)$ and $UT_jU^* = \tilde{T_j}$ for all $j=1, \ldots, d.$ This completes the proof.
\end{proof}

\begin{proposition}\label{unit-wt-eq-cor}
Let $d$ be a positive integer and $H$ be a complex Hilbert space. Let $T = (T_1, \ldots, T_d)$ be a commuting operator-valued multishift on $\ell^2_H(\mathbb N^d)$ with unitary operator weights $\{A^{(j)}_{\alpha} : \alpha \in \mathbb N^d,\  j=1, \ldots, d\}$. Then $T$ satisfies the von Neumann's inequality.
\end{proposition}

\begin{proof}
It follows from the preceding lemma that $T$ is unitarily equivalent to the operator-valued multishift on $\ell^2_H(\mathbb N^d)$ with operator weights being the identity operator on $H$. In other words, $T$ is unitarily equivalent to the unweighted multishift on $\ell^2_H(\mathbb N^d)$. Since the unweighted multishift on $\ell^2_H(\mathbb N^d)$ dilates to the (unweighted) bilateral multishift on $\ell^2_H(\mathbb Z^d)$, which is a $d$-tuple of commuting unitary operators, it follows that $T$ satisfies the von Neumann's inequality.
\end{proof}

Let us see another family of commuting operator-valued multishifts for which the von Neumann's inequality always holds.

\begin{proposition}\label{decom-classical}
Let $n,$ $d$ be positive integers and $T = (T_1, \ldots, T_d)$ be a commuting operator-valued multishift on $\ell^2_{\mathbb C^n}(\mathbb N^d)$ with operator weights given by 
\beqn
A^{(j)}_{\alpha} = \begin{pmatrix} w_{1,\alpha}^{(j)} & 0 & \ldots & 0 \\ 0 & w_{2,\alpha}^{(j)} & \ldots & 0\\
\vdots & \vdots & \ddots & \vdots\\
0 & 0& \ldots & w_{n,\alpha}^{(j)} \end{pmatrix} \mbox{ for all } \alpha \in \mathbb N^d \mbox{ and } j = 1, \ldots, d.
\eeqn
Then $T$ is unitarily equivalent to $W_1 \oplus \cdots \oplus W_n$, where $W_k$, $k = 1, \ldots, n$, is the classical multishift on $\ell^2(\mathbb N^d)$ with weights $\{w_{k,\alpha}^{(j)} : \alpha \in \mathbb N^d,\ j=1,\ldots, d\}$. 
\end{proposition}

\begin{proof}
Let $\{e_1, \ldots, e_n\}$ be the standard orthonormal set of vectors in $\mathbb C^n$. For $k=1, \ldots, n$, define 
\beqn
e_{\alpha, k} = \oplus_{\beta \in \mathbb N^d} x_\beta \mbox{ where } x_\beta = 0 \mbox{ if } \beta \neq \alpha \mbox{ and } x_\alpha = e_k.
\eeqn
For each $k = 1, \ldots n$, set 
\beqn
\mathcal M_k := \bigvee \big\{e_{\alpha, k} : \alpha \in \mathbb N^d \big\}. 
\eeqn
Then $\ell^2_{\mathbb C^n}(\mathbb N^d) = \mathcal M_1 \oplus \cdots \oplus \mathcal M_n$. Further, it is a routine verification to see that $\mathcal M_k$ is a reducing subspace of each $T_j$ and $(T_1|_{\mathcal M_k}, \ldots, T_d|_{\mathcal M_k})$ is unitarily equivalent to a classical multishift $W_k$ with weights $\{w_{k,\alpha}^{(j)} : \alpha \in \mathbb N^d,\ j=1,\ldots, d\}$ for $k=1, \ldots, n$. This completes the proof.
\end{proof}

As the von Neumann's inequality respects the direct sum, the following corollary immediately follows from Theorem \ref{Hartz} and the preceding proposition. 

\begin{corollary}
Let $n,$ $d$ be positive integers and $T = (T_1, \ldots, T_d)$ be a commuting contractive operator-valued multishift on $\ell^2_{\mathbb C^n}(\mathbb N^d)$ with operator weights being invertible $n \times n$ diagonal matrices. Then $T$ satisfies the von Neumann's inequality.
\end{corollary}

The following proposition facilitates us to produce a class of operator-valued multishifts for which the von Neumann's inequality does not hold.

\begin{proposition}\label{prop-for-c-example}
Let $H$ be a complex Hilbert space and $d$ be a positive integer. Suppose that $A = (A_1, \ldots, A_d)$ is a $d$-tuple of commuting contractions on $H$ and $T = (T_1, \ldots, T_d)$ is the commuting operator-valued multishift on $\ell^2_{H}(\mathbb N^d)$ with operator weights given by $A_\alpha^{(j)} = A_j$ for all $\alpha \in \mathbb N^d$ and $j=1, \ldots, d$. Then $T$ satisfies the von Neumann's inequality if and only if $A$ satisfies the von Neumann's inequality. 
\end{proposition}

\begin{proof}
Observe that the Hilbert space $\ell^2_{H}(\mathbb N^d)$ can be realized as $H \otimes \ell^2(\mathbb N^d)$. Hence it is not difficult to see that 
\beqn
T_j = A_j \otimes S_j \mbox{ for all } j = 1, \ldots, d,
\eeqn 
where $S = (S_1, \ldots, S_d)$ is the unweighted multishift on $\ell^2(\mathbb N^d)$. Note that $S$ is unitarily equivalent to the $d$-tuple of operators of multiplication by the coordinate functions on the Hardy space of the polydisc $\mathbb D^d$. Hence it follows that the Taylor spectrum of $S$ is $\overline{\mathbb D}^d$. Since the Taylor spectrum is contained in the algebraic spectrum, it follows that $(1, \ldots, 1)$ is in the algebraic spectrum of $S$. Now the desired conclusion is immediate from Theorem \ref{intro-thm}.
\end{proof}

\begin{corollary}
Let $d$ be a positive integer and $A = (A_1, \ldots, A_d)$ be a $d$-tuple of commuting $2 \times 2$ or $3 \times 3$ contractive matrices. Let $T = (T_1, \ldots, T_d)$ be the commuting operator-valued multishift on $\ell^2_{\mathbb C^n}(\mathbb N^d)$ $(n=2 \mbox{ or } 3)$ with operator weights given by $A_\alpha^{(j)} = A_j$ for all $\alpha \in \mathbb N^d$ and $j=1, \ldots, d$. Then $T$ satisfies the von Neumann's inequality.
\end{corollary}

\begin{proof}
The fact that any $d$-tuple of commuting $2 \times 2$ contractive matrices satisfies the von Neumann's inequality was established in \cite{D} while that for a $d$-tuple of commuting $3\times 3$ contractive matrices was proved in \cite{K}. Now rest of the proof is immediate from the preceding proposition.
\end{proof}

We are now ready to give the example which we mentioned in the beginning of this text. Our example is motivated from the one given by Kaijser and Varopoulos \cite{V1} to disprove the von Neumann's inequality for $3$-tuple of commuting contractions. Before this, observe that if a commuting $d$-tuple $V$ of contractive matrices is a counter-example to the von Neumann's inequality, then for some $t \in (0,1)$ the commuting $d$-tuple $A = tI + (1-t)V$ of invertible contractive matrices is also a counter-example to the von Neumann's inequality. Hence, if $S$ is the unweighted multishift on $\ell^2(\mathbb N^d)$, then the operator-valued multishift $A\otimes S$ satisfies all the conditions of Theorem \ref{intro-thm} and is the desired counter-example to the von Neumann's inequality.
The following example illustrates this discussion.
\begin{example}
Let $c\in (0,1/(6+\sqrt{30}))$. Following \cite[Definition 2.5]{GR} (see also \cite{V2}), consider the Varopoulos operators  on $\mathbb C^4$ given by
	\[V_j=
	\left(
	\begin{array}{cccc}
		0 & x_j & y_j& 0\\
		0 & 0 & 0 & x_j\\
		0 & 0 & 0 & y_j\\
		0 & 0 & 0 & 0\\
	\end{array}
	\right), \quad  j=1,2,3,\]
	where $x_j,y_j\in\mathbb R$ and $x_j^2+y_j^2=(1-c)^2$ for each $j=1,2,3.$ Set $A_j:=cI+V_j$ for all $j=1,2,3.$ Since $(V_1,V_2,V_3)$ is a commuting tuple, it follows that $A=(A_1,A_2,A_3)$ is also a commuting $3$-tuple of invertible matrices. Moreover, $\|A_j\|\leq c+\|V_j\|=1$ for all $j=1,2,3.$ 
	Let $T = (T_1, T_2, T_3)$ be the commuting operator-valued multishift on $\ell^2_{\mathbb C^4}(\mathbb N^d)$ with operator weights given by $A_\alpha^{(j)} = A_j$ for all $\alpha \in \mathbb N^d$ and $j=1, 2, 3$. 
	Now, consider the Varopoulos-Kaijser polynomial \cite{V1} (see also \cite{Gu})
	$$p_{\!_V}(z_1,z_2,z_3):=z_{1}^{2}+z_{2}^{2}+z_{3}^{2}-2z_1z_2-2z_2z_3-2z_3z_1.$$
	It is shown in \cite{V1} (see also \cite{Hol}, \cite{GR}) that $\sup_{z\in\mathbb D^3}|p_{\!_V}(z)|=5.$ Further, we observe that $p_{\!_V}(A_1,A_2,A_3)$ is given by
\beqn
	\left(
	\begin{smallmatrix}
	\displaystyle\sum_{j,k=1}^{3}c^2 a_{jk} &\displaystyle \sum_{j,k=1}^{3}c\, a_{jk}(x_j+x_k) & \displaystyle\sum_{j,k=1}^{3}c\, a_{jk}(y_j+y_k) & \displaystyle\sum_{j,k=1}^{3}a_{jk}\langle X_j,X_k\rangle\\
	0 & \displaystyle\sum_{j,k=1}^{3}c^2 a_{jk}  & 0 &  \displaystyle \sum_{j,k=1}^{3}c\, a_{jk}(x_j+x_k)\\
	0 & 0 &   \displaystyle\sum_{j,k=1}^{3}c^2 a_{jk}  & \displaystyle\sum_{j,k=1}^{3}c\, a_{jk}(y_j+y_k)\\
	0 & 0 & 0 & \displaystyle\sum_{j,k=1}^{3}c^2 a_{jk}
	\end{smallmatrix}
	\right),
	\eeqn	
	where \begin{equation*}
	(a_{jk}):=
	\left(
	\begin{array}{rrr}
		1 & -1 & -1\\
		-1 & 1 & -1\\
		-1 & -1 & 1\\
	\end{array}
	\right),
\end{equation*}
and $X_j=(x_j,y_j)$ for $j=1,2,3.$
	From above it can be concluded that 
\beqn
\|p_{\!_V}(A_1,A_2,A_3)\|&\Ge& \Big|\sum_{j,k=1}^{3}a_{jk}\langle X_j,X_k\rangle\Big|\\
&=& 3(1-c)^2-2(\langle X_1,X_2\rangle + \langle X_2,X_3\rangle + \langle X_3,X_1\rangle).
\eeqn	
Following \cite[Lemma 2.18]{Thesis}, it can be shown that the right hand side of the above inequality achieves its maximum value at $X_1=(1-c)(1,0),\, X_2=(1-c)(-1/2,\sqrt{3}/2)$ and $X_3=(1-c)(-1/2,-\sqrt{3}/2)$ and therefore
\beqn
\|p_{\!_V}(A_1,A_2,A_3)\|\Ge 6(1-c)^2.
\eeqn
Thus using this and the facts that $c<1/(6+\sqrt{30})$ and $\sup_{z\in\mathbb D^3}|p_{\!_V}(z)|=5$, we conclude that $\|p_{\!_V}(A_1,A_2,A_3)\|>\sup_{z\in\mathbb D^3}|p_{\!_V}(z)|$. Now from Proposition \ref{prop-for-c-example}, we deduce that the operator-valued multishift $T=(T_1,T_2,T_3)$ does not satisfy the von Neumann's inequality.\end{example}

\medskip \textit{Acknowledgment}.
The authors are grateful to Gadadhar Misra and Sameer Chavan for several helpful suggestions and constant support. We express our gratitude to the faculty and the administration of Department of Mathematics and Statistics, IIT Kanpur and Department of Mathematics, IISc Bangalore for their warm hospitality during the preparation of this paper. We sincerely thank the referee for several constructive comments which considerably improved the presentation of this paper.

\end{document}